OpenStudy (anonymous):

Let p>=3 and G be a planar graph with p vertices and 3p-7 edges. Prove G is connected. A connected graph is a graph in which for any 2 vertices there is a path connecting them.

OpenStudy (anonymous):

A graph is planar if you can draw it in a plane so that edges only intersect at their ends and no two vertices coincide

OpenStudy (jtvatsim):

Perhaps it would be possible to argue by mathematical induction...

OpenStudy (jtvatsim):

The first case is p = 3. Then, we have 3 vertices and 3(3) - 7 = 2 edges. It is fairly obvious that G is connected in this case.

OpenStudy (anonymous):

I'm not especially good with proofs by induction. I can see that the base case is true however.

OpenStudy (jtvatsim):

Yeah, idk... the induction hypothesis would be suppose that G is connected for p = k with 3k - 7 edges.

OpenStudy (jtvatsim):

Then, for the (k+1)st case, we would have k+1 vertices and 3(k+1)-7 = 3k-7 + 3 edges.

OpenStudy (jtvatsim):

We claim that G is connected. Observe that the first k vertices can be connected using the first (3k-7) edges by the induction hypothesis. So this "subgraph" is connected. Then, we have 1 vertex left and 3 edges remaining. But we can easily attached this 1 vertex to the connected "subgraph" with the 3 edges. Hence, p = k+1 is also connected.

OpenStudy (jtvatsim):

It seems to work, but I'll be the first to admit that I don't know much graph theory. :P

OpenStudy (jtvatsim):

I think the statement that "we can easily attach this 1 vertex to the connected 'subgraph'" is probably a bit fuzzy. This probably has something to do with a planar property.

OpenStudy (jtvatsim):

A different approach might be to use the "evenness" or "oddness" of the vertices. I'm literally just throwing ideas out here.

OpenStudy (anonymous):

That seems like it could work, but I think the problem is whether or not adding the edges would destroy the planarity

OpenStudy (jtvatsim):

Yeah, but the thing to keep in mind is that the planarity is given... it's the connectivity that is the concern.

OpenStudy (jtvatsim):

Not sure, this one's got me. Good luck!

OpenStudy (anonymous):

Thanks for your help anyways!